# The Impact of Domain Knowledge and Multi-Modality on Intelligent Molecular Property Prediction: A Systematic Survey

### v1

@@ 4.1.3 Transformer-based
Originally excelling in NLP, the Transformer architecture is renowned for its self-attention mechanism, which allows for parallel processing of entire sequences. This capability enables it to efficiently manage long-range dependencies within data, making it highly effective in MPP. Its adeptness at understanding detailed contextual relationships enhances the accuracy and computational efficiency in predictive modeling. Wang et al.[87] and Chithrananda et al.[67] use Transformer to extract molecular information from SMILES, which is treated as natural language. Wang et al.[88] proposed two significant advances in molecular data processing: structural fingerprint tokenization for more efficient molecule graph tokenization and normalized graph raw shortcut-connection to enhance latent representations in complex model structures. To address challenges in the validity and robustness of SMILES representations, Yüksel et al.[68] uniquely utilizes SELFIES, a robust and flexible molecular representation format, to learn high-quality molecular features, enhancing the reliability of molecular data analysis in computational chemistry. To predicting activity coefficients in binary mixtures, Winter et al.[89] integrate information from two SMILES strings representing the mixture components, along with temperature and token position data, into a unified matrix for input encoding. Ross et al.[12] explore the diference between absolute and relative position embeddings in representing molecular SMILES and provide a linear attention approximation of the recently proposed relative position RoFormer[90]. Ross et al.[12] delves into the differences between absolute and relative position embeddings in SMILES representation, proposing an efficient linear attention approximation for the RoFormer[90] model, which focuses on relative positioning, to enhance molecular SMILES processing in deep learning applications.
@@
The Transformer architecture, originally designed for sequence data, has been effectively adapted for molecular graph representation in recent research. Its proficiency in handling global molecular information enhances its utility in molecular property prediction, showcasing its versatility beyond traditional sequence analysis. Maziarka et al.[91] proposed the molecule attention transformer, which adapts the Transformer architecture, augmenting the self-attention mechanism with inter-atomic distances and molecular graph structure. Li et al.[9] focus on chemical bonds in molecular representations, employing molecular line graphs to illustrate edge adjacencies in original molecular graphs. Each graph is augmented with a knowledge node containing molecular descriptors and fingerprints, connected to its original nodes Rong et al.[11] combines message passing networks with a transformer-style architecture, extract vectors as queries, keys and values from nodes of the graph, then feed them into the attention block. Yin et al.[13] developed a method that alternates between GNN and Transformer layers, repeated in sequence. This approach effectively blends local and global information, allowing the graph transformer to comprehensively integrate node data from both nearby and distant sources. To extrace the coarse-grained view, Ren et al.[92] make the molecular graph first enters the message passing phase of the traditional GNN layers to update the node embeddings, then enters graph transformation layers to learn different granular information. To achieve one encoder for extracting 2D or 3D information, Luo et al.[14] use two separated channels to encode 2D and 3D structural information and incorporate them with the atom features in the network modules. To fully leverages chemical knowledge, Gao et al.[93] construct an embedding unit comprising a GNN and a transformer to balance the neighboring and distant interactions of an atom, and more attention is given to conjugated systems, unsaturated bonds, heteroatoms and the molecular topology. To extract the molecular fragment information, Jiang et al.[94] design a pharmacophoric-constrained multi-views molecular representation graph, enabling PharmHGT to extract vital chemical information from functional substructures and chemical reactions.
@@ 4.1.5 Multi-modality
In conclusion, the above work underscores the effectiveness of multi-modal learning in the context of MPP. This approach facilitates the seamless integration of various molecular data formats, including sequences, graph data types, and molecular descriptors. By amalgamating these diverse sources of information, multi-modal learning provides a richer and more nuanced understanding of molecular properties, which is essential for achieving accurate predictions.
@@ 4.2 Training Strategy
In this section, we introduce all approaches used to train DL models. While supervised learning has been traditionally predominant, its reliance on scarce labeled data presents limitations.
Figure 6: Training strategy: The high cost of experiments often results in a scarcity of labeled data, leading to challenges like overfitting and poor generalization. To address this, our figure illustrates various advanced training strategies. a). Self-supervised learning, such as contrastive learning and masking recovery, utilizes unlabeled data for pretraining encoders, enabling them to learn molecular domain knowledge effectively. b). Semi-supervised learning methods like pseudo-labeling and co-training leverage both labeled and unlabeled data, enhancing the encoder’s understanding of data distribution. c). Transfer learning strategy capitalizes on pre-trained models from source tasks to boost performance on target tasks. d). Multi-task learning approaches combine related datasets to predict multiple properties simultaneously, benefiting from the relational aspects of molecular data.
To circumvent this, recent approaches have shifted towards unsupervised, self-supervised, and semi-supervised learning methods, capitalizing on the abundance of unlabeled data. Transfer learning is also employed to utilize models pre-trained on unrelated data, enhancing the model’s performance on specific tasks. Additionally, multi-task learning strategies are adopted to leverage related labeled data, further refining the model’s accuracy in predicting molecular properties. As Figure 6 shown, the details of training strategy are as followings.
4.2.1 Self-supervised Learning
Inspired by NLP, Wang et al.[87], Chithrananda et al.[67], Zhang et al.[112], Ahmad et al.[113] and Irwin et al.[114] employ masked language modeling(MLM) on large scale unlabeled data to generate context-sensitive representation, treating SMILES as natural language. Ma et al.[3] use auto-encoder strategy in pretrain stage, first convert SMILES to a vector representation and then reconstructed representation back to SMILES to update the network. Furthermore, Guo et al.[21] fusion the molecular graph and SMILES representation to recontruct the SMILES. Except using SMILES as input, Yuksel et al.[68] employs MLM in SELFIES representations in order to obtain their concise, flexible, and meaningful representations. To let nodes appearing in similar structural contexts to nearby embeddings, Hu et al.[115] propose a context prediction task by using subgraphs to predict their surrounding graph structures.
@@ 4.2.1 Self-supervised Learning
Figure 7: The training strategy summary.
@@
To address GNN oversmoothing and encourage latent node diversity, Godwin et al.[126] employ denoise technique in which they corrupt the input graph with noise, and add a noise correcting node-level loss. Zeng et al.[15] implemented an auto-encoder for molecular image reconstruction, using a discriminator to distinguish between real and fake molecular images. To expand atom vocabulary, Xia et al.[120] use a context-aware tokenizer to encode atom attributes into meaningful discrete codes, then randomly masking and recovering these codes to efficiently pretrain their encoder. Intrinsically, for molecules, a more natural representation is based on their 3D geometric structures, which largely determine the corresponding physical and chemical properties. To overcome the challenge of attaining the coordinate denoising objective, Liu et al.[127] employ an SE(3)-invariant score matching strategy to successfully transform such objective into the denoising of pairwise atomic distances. To capture the anisotropic characteristic of molecules, Feng et al.[128] propose a novel hybrid noise strategy, including noises on both dihedral angel and coordinate, and also decouple the two types of noise and design a novel fractional denoising method, which only denoises the latter coordinate part. For effectively learning 3D spatial representation, Zhou et al.[66] employ 3D position recovery and masked atom prediction as pretrain task. Further more, Jiao et al.[129] exploit the Riemann-Gaussian distribution to ensure the loss to be E(3)-invariant, enabling more robustness. To guild by the molecular domain knowledge and extract chemical information like chemists, Li et al.[9, 24] leverages the molecular descriptors and fingerprints, which serves as the semantics lost in the masked graph to guide the prediction of the masked nodes, thus making the model capture the abundant structural and semantic information from large-scale unlabeled molecules. Wu et al.[121] proposed atom property prediction to discern finer differences between atoms, and MACCS fingerprints prediction, enabling their model to extract and learn predefined molecular features. Gao et al.[130] use atom charges and 3D geometries as inputs, with molecular energies as the target labels, aiming to effectively leverage energy information for enhanced molecular analysis. To optimize multi-task integration and avoid ineffective transfer, Wang et al.[wang2023automated] introduce a fusion strategy that utilizes a surrogate metric based on the total energy of all atoms in a molecule during the pre-training stage. Zang et al.[73] designs three generative tasks that predict bond links, atom types, and bond types with the atom representations and designs two predictive tasks that predict the number of atoms and bonds with the molecule representation. Zeng et al.[131] and Broberg et al.[132] have developed methods to predict the product molecular SMILES based on the reactant molecular embedding. This approach allows for the extraction of chemical information from chemical reactions, providing insights into the molecular transformations involved in the reaction process.
@@
Contrastive learning, a method distinguishing positive and negative molecule pairs, has become a key strategy in encoder pretraining for its ability to enhance molecular structure discernment. This technique is extensively utilized in numerous studies, making it a cornerstone for improving molecular structure recognition in various models. For the SMILES augmentation, Wu et al.[52] and Zhang et al.[133] implemented SMILES enumeration, a technique that varies starting atoms and traversal orders to represent a molecule with different SMILES, thereby uncovering more intricate patterns from complex SMILES structures. Wu et al.[121] and Abdel et al.[134] utilized SMILES permutation as a data augmentation technique, involving the rearrangement of atoms in a SMILES string to create different representations without altering the underlying molecular structure. For molecular graph augmentation, techniques like node dropping, edge perturbation, attribute masking, and subgraph masking are commonly used[119, 22, 135, 136]. However, these random masking methods may not effectively guide the encoder to identify the most crucial chemical information, and might result in the creation of less accurate positive and negative molecule pairs for the training process. To capture important molecular structure and higher order semantic information, Liu et al.[137] adopted the graph attention network as the molecular graph encoder, and leveraged the learned attention weights as masking guidance to generate molecular augmentation graphs. Lin et al.[138] first models the underlying semantic structure of the graph data via clustering semantically similar graphs to select the positive and negative pair and then reweights its negative samples based on the distance between their prototypes and the query prototype such that those negatives having moderate prototype distance enjoy relatively large weights. Cui et al.[139] utilize the GNN encoder and its momentum-update version[140] to generate positive samples at the representation level, and select the negative pairs by the semantic importance of nodes, which is calculated by eigenvector centrality iteration[141]. Wang et al.[142] employ a generative probabilistic model to learn molecular graph structures for topology augmentations and simultaneously develop feature selectors to mask less critical atom features, thus generating effective attribute-level augmentations. To gain deeper insights into chemical information, many researchers incorporate domain knowledge into their contrastive learning approaches. By using backbone and side-chain information, Liu et al.[143] employ side-chain repetition, side-chain generation, backbone disruption, and backbone disruption + side-chain deletion strategy to generate hard positive, soft positive, soft negative and hard negative samples, respectively. Sun et al.[117] replaced a valid substructure by a bioisostere that introduces variation without altering the molecular properties too much, and treat them as positive pairs. Also they optimize the similarity of molecule pairs embedding to be close to the similarity of their ECFP. To avoid faulty negative pairs, Wang et al.[118] mitigate negative contrastive instances by considering ECFP similarities between molecule pairs. Wang et al.[144] calculate the weight vector using the self-attention mechanism to determine the selection probability of each character in SMILES and generate positive samples using three masked strategies: roulette masking, top masking, and random masking. To maintain semantics between conformers, Moon et al.[145] randomly selects molecules from the conformer pool instead of selecting the most stable molecules to learn the 3D structure abundantly. Kuang te al.[146] consider conformations with the same SMILES as positive pairs and the opposites as negative pairs, while keeping the weight to indicate the 3D conformation descriptor and fingerprint similarity. Knowledge graph (KG) is a semantic network composed of entities and their relations in the real world.[147] Hua et al.[148] use the atoms in SMILES as indices to query the embedding matrix to obtain entity and relation embeddings. For the entity and relation vectors of different atoms, they obtain the entity and relation embeddings of the SMILES through linear mapping, and finally concatenate the two vectors to obtain the final embedding representation. Fang et al.[149] first construct a Chemical Element KG based on periodic table of elements, to describes the relations between elements and their basic chemical attributes, Furthermore, they[23] construct another chemical Element KG based on the periodic table and Wikipedia pages to summarize the basic knowledge of elements and the closely related substructure. Those KG offers a comprehensive and standardized view from a chemical element perspective, and help to augment the original molecular graph with the guidance of KG.
@@
In the realm of MPP, a deep understanding of molecular substructures is increasingly recognized as crucial. Many recent studies leverage this domain knowledge to effectively identify and analyze important substructural information, significantly enhancing the understanding of molecular behavior. Xu et al.[150] aimed to preserve local similarities between graph instances by aligning embeddings of related subgraphs and differentiating these from unrelated pairs. They also implemented hierarchical prototypes to represent the latent distribution of graph datasets, enhancing data likelihood with respect to both GNN parameters and these hierarchical structures. Xu et al.[151] aim to capture both intra- and inter-space relationships in group representations and also introduced an attention-based mechanism for generating groups, aimed at identifying diverse molecular substructures effectively. Wang et al.[118] employ BRICS to decompose different substructures which are considered as contrastive negative pairs. Shen et al.[152] and Rong et al.[11] use fingerprints as self-generated labels, Given a graph, its network motifs like chemical functional groups or fragments, which are self-labeled by whether or not they exist in the graph. To learn the local semantics, Luo et al.[153] use graph clustering techniques to partition each whole graph into several subgraphs while preserving as much semantic information as possible, and treat the molecular graph and the clustering graph as postive pair. Benjamin et al.[154] extract substructure information by setting the junction tree(through a tree decomposition algorithm) reconstruction and fingerprint prediction task. To analyze molecular GNN strictly in terms of chemically meaningful fragments, Wu et al.[122] identifies the most crucial set of substructures(BRICS, Murckoand functional groups) in a molecule that are responsible for a model’s prediction. HeGCL[155] introduce the meta-path view that provides semantic information, and encodes graph embeddings by maximizing mutual information between global and semantic representations obtained from the outline and meta-path view, respectively. Hierarchical Molecular Graph is a usually way to extract the substructure molecular representations. Zhu et al.[77] extracts hierarchical information by utilizing co-representation learning of molecular graphs and chemically synthesizable BRICS fragments, and also uses a feature-wise attention block to adaptively recalibrate atomic features after the message passing phase. Kim et al.[123] construct a bag of fragments from a molecule through fragmentation, treating a complete or incomplete bag as a positive or negative view of the original molecule, respectively. Xie et al.[156] proposed a fragment-based molecular graph (FMG) to represent the topological relationship between chemistry-aware substructures within a molecule. They then pre-train it on a fragment level using contrastive learning with well-designed hard negative pairs to extract node representations in FMGs. Xie et al.[156] employ contrastive learning to train fragment encoder, and treat a fragment graph and its associated SMILES string as positive pair, also treat an extremely similar fragment graph serves as a hard negative sample generated by a hard negative construction module. And then encoder the fragment-based molecular graph to extract molecular representation. Xie et al.[156] used contrastive learning for fragment encoder training, pairing a fragment graph with its SMILES string as positive, and pairing a similar fragment graph as negative which generated by a negative construction module, and then applied to fragment-based molecular graphs, enhances molecular representation extraction. Ji et al.[157] decompose the molecular graph using a more reasonable method to construct the fragment graph. They select positive/negative pairs based on similarities between two-level molecule pairs and employ a contrastive loss function, as proposed by Hadsell et al., to pretrain the encoder.
@@
Diverse data formats have been shown to be crucial for MPP, and the multi-modal approach, merging these formats, enhances prediction accuracy by offering a holistic view of molecules. This technique, increasingly adopted in research, combines different data types for a more detailed molecular analysis. To leverage two popular molecular representations and augmentations for each modality, Pinheiro et al.[158], Zhang et al.[159], Zhu et al.[116] and Sun et al.[102] exploit two molecular representations that can be easily acquired from chemical space: the SMILES string and the molecular graph, and then make them as positive pairs. Li et al.[19] utilized self-supervised learning by exploiting the relationship and consistency between 2D topological and 3D geometric structures of molecules. Additionally, Liu et al.[18] applied a generative self-supervised learning approach that focuses on intra-data knowledge, reconstructing key features at the individual data point level to enhance the understanding of molecular structures. 3D Infomax[160] maximized the mutual information between learned 3D summary vectors and the representations of a GNN. Zhu et al.[20] implemented a multifaceted pretraining strategy involving the reconstruction of masked atoms and coordinates, generating 3D conformations based on 2D graphs, and creating 2D graphs from 3D conformations. Kim et al.[123] focused on extracting explicit 3D geometric information by proposing a solution for predicting torsional angles between adjacent molecular fragments, thereby enhancing the depth and accuracy of 3D molecular analysis. Zhu et al.[161] aimed to integrate multiple molecular feature views, including 2D and 3D graphs, Morgan fingerprints, and SMILES strings, ensuring cohesive embedding consistency between these diverse representations for a more unified molecular analysis. The reviewed works show that self-supervised learning, particularly through methods like encoder-recovery and contrastive learning, effectively utilizes unlabeled data to improve model generalization in MPP. These methods excel in learning prior knowledge through various pretraining tasks, allowing for integration of multi-modal data and domain knowledge. This approach significantly enhances the adaptability and performance of models in molecular property prediction scenarios.

### v3

@@ 4.1.3 Transformer-based
Originally excelling in NLP, the Transformer architecture is renowned for its self-attention mechanism, which allows for parallel processing of entire sequences. This capability enables it to efficiently manage long-range dependencies within data, making it highly effective in MPP. Its adeptness at understanding detailed contextual relationships enhances the accuracy and computational efficiency in predictive modeling.
@@
Wang et al.[87] and Chithrananda et al.[55] use Transformer to extract molecular information from SMILES, which is treated as natural language. Wang et al.[88] proposed two significant advances in molecular data processing: structural fingerprint tokenization for more efficient molecule graph tokenization and normalized graph raw shortcut-connection to enhance latent representations in complex model structures. To address challenges in the validity and robustness of SMILES representations, Yüksel et al.[56] uniquely utilizes SELFIES, a robust and flexible molecular representation format, to learn high-quality molecular features, enhancing the reliability of molecular data analysis in computational chemistry. To predicting activity coefficients in binary mixtures, Winter et al.[89] integrate information from two SMILES strings representing the mixture components, along with temperature and token position data, into a unified matrix for input encoding. Ross et al.[12] delved into the differences between absolute and relative position embeddings in SMILES representation, proposing an efficient linear attention approximation for the RoFormer[90] model, which focuses on relative positioning, to enhance molecular SMILES processing in deep learning applications.
@@
The Transformer architecture, originally designed for sequence data, has been effectively adapted for molecular graph representation in recent research. Its proficiency in handling global molecular information enhances its utility in molecular property prediction, showcasing its versatility beyond traditional sequence analysis. Maziarka et al.[91] proposed the Molecule Attention Transformer, which adapts the Transformer architecture, augmenting the self-attention mechanism with inter-atomic distances and molecular graph structure. Li et al.[9] focus on chemical bonds in molecular representations, employing molecular line graphs to illustrate edge adjacencies in original molecular graphs. Each graph is augmented with a knowledge node containing molecular descriptors and fingerprints, connected to its original nodes Rong et al.[11] combines message passing networks with a Transformer-style architecture, extract vectors as queries, keys and values from nodes of the graph, then feed them into the attention block. Park et al.[92] introduced Graph Relative Positional Encoding, which effectively encodes graph structures by concurrently addressing node-topology and node-edge interactions, bypassing the need for linearization. Hussain et al.[93] developed the Edge-augmented Graph Transformer, employing global self-attention rather than traditional static convolutional aggregation. This design facilitates dynamic, long-range node interactions and incorporates edge channels for evolving structural information, enabling direct predictions on edges and links. Masters et al.[94] integrate a substantial message-passing module with a biased self-attention layer to facilitate both localized biases and broad-scale communication. Chen et al.[95] proposed Graph Propagation Attention, which explicitly handles node-to-node, node-to-edge, and edge-to-node interactions, allowing for comprehensive information propagation. Yin et al.[13] developed a method that alternates between GNN and Transformer layers, repeated in sequence. This approach effectively blends local and global information, allowing the Graph Transformer to comprehensively integrate node data from both nearby and distant sources. To extrace the coarse-grained view, Ren et al.[96] make the molecular graph first enters the message passing phase of the traditional GNN layers to update the node embeddings, then enters graph transformation layers to learn different granular information. To achieve one encoder for extracting 2D or 3D information, Luo et al.[14] use two separated channels to encode 2D and 3D structural information and incorporate them with the atom features in the network modules. To fully leverages chemical knowledge, Gao et al.[97] construct an embedding unit comprising a GNN and a Transformer to balance the neighboring and distant interactions of an atom, and more attention is given to conjugated systems, unsaturated bonds, heteroatoms and the molecular topology. To extract the molecular fragment information, Jiang et al.[98] design a pharmacophoric-constrained multi-views molecular representation graph, enabling PharmHGT to extract vital chemical information from functional substructures and chemical reactions.
@@ 4.1.5 Multi-modality-based
In conclusion, the above work underscores the effectiveness of multi-modal learning in the context of MPP. This approach facilitates the seamless integration of various molecular modality, including sequences, graph data types, and molecular descriptors. By amalgamating these diverse sources of information, multi-modal learning provides a richer and more nuanced understanding of molecular properties, which is essential for achieving accurate predictions.
@@ 4.2 Training Strategy
In this section, we introduce all approaches used to train DL models. While supervised learning has been traditionally predominant, its reliance on scarce labeled data presents limitations. To circumvent this, recent approaches have shifted towards unsupervised, self-supervised, and semi-supervised learning methods, capitalizing on the abundance of unlabeled data. Transfer learning is also employed to utilize models pretrained on unrelated data, enhancing the model’s performance on specific tasks. Additionally, multi-task learning strategies are adopted to leverage related labeled data, further refining the model’s accuracy in predicting molecular properties. As Figure 6 and Figure 7 shown, the details of training strategy are as followings.
@@
Figure 6: The training strategy summary. We categorize training strategies into four key types: Self-Supervised Learning, Semi-Supervised Learning, Transfer Learning, and Multi-Task Learning. Each category includes a detailed description of the main focuses and considerations prevalent in renowned studies, illustrating the diverse approaches and priorities within each training strategy for optimizing molecular property prediction.
4.2.1 Self-supervised Learning
@@
Inspired by NLP, Wang et al.[87], Chithrananda et al.[55], Zhang et al.[126], Ahmad et al.[127] and Irwin et al.[128] employ masked language modeling(MLM) on large scale unlabeled data to generate context-sensitive representation, treating SMILES as natural language. Ma et al.[3] use auto-encoder strategy in pretrain stage, first convert SMILES to a vector representation and then reconstructed representation back to SMILES to update the network. Furthermore, Guo et al.[21] fusion the molecular graph and SMILES representation to recontruct the SMILES. Except using SMILES as input, Yuksel et al.[56] employs MLM in SELFIES representations in order to obtain their concise, flexible, and meaningful representations. To let nodes appearing in similar structural contexts to nearby embeddings, Hu et al.[129] propose a context prediction task by using subgraphs to predict their surrounding graph structures. To address GNN oversmoothing and encourage latent node diversity, Godwin et al.[130] employ denoise technique in which they corrupt the input graph with noise, and add a noise correcting node-level loss. Zeng et al.[15] implemented an auto-encoder for molecular image reconstruction, using a discriminator to distinguish between real and fake molecular images. To expand atom vocabulary, Xia et al.[118] use a context-aware tokenizer to encode atom attributes into meaningful discrete codes, then randomly masking and recovering these codes to efficiently pretrain their encoder. Intrinsically, for molecules, a more natural representation is based on their 3D geometric structures, which largely determine the corresponding physical and chemical properties. To overcome the challenge of attaining the coordinate denoising objective, Liu et al.[131] employ an SE(3)-invariant score matching strategy to successfully transform such objective into the denoising of pairwise atomic distances. To capture the anisotropic characteristic of molecules, Feng et al.[132] propose a novel hybrid noise strategy, including noises on both dihedral angel and coordinate, and also decouple the two types of noise and design a novel fractional denoising method, which only denoises the latter coordinate part. For effectively learning 3D spatial representation, Zhou et al.[54] employ 3D position recovery and masked atom prediction as pretrain task. Further more, Jiao et al.[133] exploit the Riemann-Gaussian distribution to ensure the loss to be E(3)-invariant, enabling more robustness. To guild by the molecular domain knowledge and extract chemical information like chemists, Li et al.[9, 24] leverages the molecular descriptors and fingerprints, which serves as the semantics lost in the masked graph to guide the prediction of the masked nodes, thus making the model capture the abundant structural and semantic information from large-scale unlabeled molecules. Wu et al.[119] proposed atom property prediction to discern finer differences between atoms, and MACCS fingerprints prediction, enabling their model to extract and learn predefined molecular features. Gao et al.[134] use atom charges and 3D geometries as inputs, with molecular energies as the target labels, aiming to effectively leverage energy information for enhanced molecular analysis. To optimize multi-task integration and avoid ineffective transfer, Wang et al.[135] introduce a fusion strategy that utilizes a surrogate metric based on the total energy of all atoms in a molecule during the pretraining stage. Zang et al.[73] designs three generative tasks that predict bond links, atom types, and bond types with the atom representations and designs two predictive tasks that predict the number of atoms and bonds with the molecule representation. Zeng et al.[136] and Broberg et al.[137] have developed methods to predict the product molecular SMILES based on the reactant molecular embedding. This approach allows for the extraction of chemical information from chemical reactions, providing insights into the molecular transformations involved in the reaction process.
@@
Contrastive learning, a method distinguishing positive and negative molecule pairs, has become a key strategy in encoder pretraining for its ability to enhance molecular structure discernment. This technique is extensively utilized in numerous studies, making it a cornerstone for improving molecular structure recognition in various models. For the SMILES augmentation, Wu et al.[50] and Zhang et al.[138] implemented SMILES enumeration, a technique that varies starting atoms and traversal orders to represent a molecule with different SMILES, thereby uncovering more intricate patterns from complex SMILES structures. Wu et al.[119] and Abdel et al.[139] utilized SMILES permutation as a data augmentation technique, involving the rearrangement of atoms in a SMILES string to create different representations without altering the underlying molecular structure. For molecular graph augmentation, techniques like node dropping, edge perturbation, attribute masking, and subgraph masking are commonly used[117, 22, 140, 141]. However, these random masking methods may not effectively guide the encoder to identify the most crucial chemical information, and might result in the creation of less accurate positive and negative molecule pairs for the training process. To capture important molecular structure and higher order semantic information, Liu et al.[142] adopted the graph attention network as the molecular graph encoder, and leveraged the learned attention weights as masking guidance to generate molecular augmentation graphs. Lin et al.[143] first models the underlying semantic structure of the graph data via clustering semantically similar graphs to select the positive and negative pair and then reweights its negative samples based on the distance between their prototypes and the query prototype such that those negatives having moderate prototype distance enjoy relatively large weights. Cui et al.[144] utilize the GNN encoder and its momentum-update version[145] to generate positive samples at the representation level, and select the negative pairs by the semantic importance of nodes, which is calculated by eigenvector centrality iteration[146]. Wang et al.[147] employ a generative probabilistic model to learn molecular graph structures for topology augmentations and simultaneously develop feature selectors to mask less critical atom features, thus generating effective attribute-level augmentations. To gain deeper insights into chemical information, many researchers incorporate domain knowledge into their contrastive learning approaches. By using backbone and side-chain information, Liu et al.[148] employ side-chain repetition, side-chain generation, backbone disruption, and backbone disruption + side-chain deletion strategy to generate hard positive, soft positive, soft negative and hard negative samples, respectively. Sun et al.[115] replaced a valid substructure by a bioisostere that introduces variation without altering the molecular properties too much, and treats them as positive pairs. Also, they optimize the similarity of molecule pairs embedding to be close to the similarity of their ECFP. To avoid faulty negative pairs, Wang et al.[116] mitigate negative contrastive instances by considering ECFP similarities between molecule pairs. Wang et al.[149] calculate the weight vector using the self-attention mechanism to determine the selection probability of each character in SMILES and generate positive samples using three masked strategies: roulette masking, top masking, and random masking. To maintain semantics between conformers, Moon et al.[150] randomly selects molecules from the conformer pool instead of selecting the most stable molecules to learn the 3D structure abundantly. Kuang te al.[151] consider conformations with the same SMILES as positive pairs and the opposites as negative pairs, while keeping the weight to indicate the 3D conformation descriptor and fingerprint similarity. Knowledge graph (KG) is a semantic network composed of entities and their relations in the real world.[152] Hua et al.[153] use the atoms in SMILES as indices to query the embedding matrix to obtain entity and relation embeddings. For the entity and relation vectors of different atoms, they obtain the entity and relation embeddings of the SMILES through linear mapping, and finally concatenate the two vectors to obtain the final embedding representation. Fang et al.[154] first construct a Chemical Element KG based on periodic table of elements, to describes the relations between elements and their basic chemical attributes, Furthermore, they[23] construct another chemical Element KG based on the periodic table and Wikipedia pages to summarize the basic knowledge of elements and the closely related substructure. Those KG offers a comprehensive and standardized view from a chemical element perspective, and help to augment the original molecular graph with the guidance of KG.
In the realm of MPP, a deep understanding of molecular substructures is increasingly recognized as crucial. Many recent studies leverage this domain knowledge to effectively identify and analyze important substructural information, significantly enhancing the understanding of molecular behavior. Xu et al.[155] aimed to preserve local similarities between graph instances by aligning embeddings of related subgraphs and differentiating these from unrelated pairs. They also implemented hierarchical prototypes to represent the latent distribution of graph datasets, enhancing data likelihood with respect to both GNN parameters and these hierarchical structures.
Figure 7: Training strategy: The high cost of experiments often results in a scarcity of labeled data, leading to challenges like overfitting and poor generalization. To address this, our figure illustrates various advanced training strategies. a). Self-supervised learning, such as contrastive learning and masking recovery, utilizes unlabeled data for pretraining encoders, enabling them to learn molecular domain knowledge effectively. b). Semi-supervised learning methods like pseudo-labeling and co-training leverage both labeled and unlabeled data, enhancing the encoder’s understanding of data distribution. c). Transfer learning strategy capitalizes on pretrained models from source tasks to boost performance on target tasks. d). Multi-task learning approaches combine related datasets to predict multiple properties simultaneously, benefiting from the relational aspects of molecular data.
@@
Wang et al.[116] employ BRICS to decompose different substructures which are considered as contrastive negative pairs. Motifs, including chemical functional groups or fragments, serve as self-generated labels determined by their presence or absence in the graph. Shen et al.[156] and Rong et al.[11] used these labels to pretrain their encoder. To learn the local semantics, Luo et al.[157] use graph clustering techniques to partition each whole graph into several subgraphs while preserving as much semantic information as possible, and treat the molecular graph and the clustering graph as postive pair. Benjamin et al.[158] extract substructure information by setting the junction tree(through a tree decomposition algorithm) reconstruction and fingerprint prediction task. To analyze molecular GNN strictly in terms of chemically meaningful fragments, Wu et al.[120] identifies the most crucial set of substructures(BRICS and Murckoand functional groups) in a molecule that are responsible for a model’s prediction. HeGCL[159] introduce the meta-path view that provides semantic information, and encodes graph embeddings by maximizing mutual information between global and semantic representations obtained from the outline and meta-path view, respectively. Hierarchical Molecular Graph is a usually way to extract the substructure molecular representations. Zhu et al.[77] extracts hierarchical information by utilizing co-representation learning of molecular graphs and chemically synthesizable BRICS fragments, and also uses a feature-wise attention block to adaptively recalibrate atomic features after the message passing phase. Kim et al.[121] construct a bag of fragments from a molecule through fragmentation, treating a complete or incomplete bag as a positive or negative view of the original molecule, respectively. Xie et al.[160] proposed a fragment-based molecular graph (FMG) to represent the topological relationship between chemistry-aware substructures within a molecule. They then pretrained it on a fragment level using contrastive learning with well-designed hard negative pairs to extract node representations in FMGs. Ji et al.[161] decompose the molecular graph using a more reasonable method to construct the fragment graph. They select positive/negative pairs based on similarities between two-level molecule pairs and employ a contrastive loss function, as proposed by Hadsell et al.[162], to pretrain the encoder.
@@
Diverse data formats have been shown to be crucial for MPP, and the multi-modal approach, merging these formats, enhances prediction accuracy by offering a holistic view of molecules. This technique, increasingly adopted in research, combines different data types for a more detailed molecular analysis. To leverage two popular molecular representations and augmentations for each modality, Pinheiro et al.[163], Zhang et al.[164], Zhu et al.[114] and Sun et al.[106] exploit two molecular representations that can be easily acquired from chemical space: the SMILES string and the molecular graph, and then make them as positive pairs. Li et al.[19] utilized self-supervised learning by exploiting the relationship and consistency between 2D topological and 3D geometric structures of molecules. Additionally, Liu et al.[18] applied a generative self-supervised learning approach that focuses on intra-data knowledge, reconstructing key features at the individual data point level to enhance the understanding of molecular structures. 3D Infomax[165] maximized the mutual information between learned 3D summary vectors and the representations of a GNN. Zhu et al.[20] implemented a multifaceted pretraining strategy involving the reconstruction of masked atoms and coordinates, generating 3D conformations based on 2D graphs, and creating 2D graphs from 3D conformations. Kim et al.[121] focused on extracting explicit 3D geometric information by proposing a solution for predicting torsional angles between adjacent molecular fragments, thereby enhancing the depth and accuracy of 3D molecular analysis. Zhu et al.[166] aimed to integrate multiple molecular feature views, including 2D and 3D graphs, Morgan fingerprints, and SMILES strings, ensuring cohesive embedding consistency between these among representations for a more unified molecular analysis.
The reviewed works show that self-supervised learning, particularly through methods like encoder-recovery and contrastive learning, effectively utilizes unlabeled data to improve model generalization in MPP. These methods excel in learning prior knowledge through various pretraining tasks, allowing for integration of multi-modal data and domain knowledge. This approach significantly enhances the adaptability and performance of models in molecular property prediction scenarios.